QUESTION 10

10.1 Use 9.2 to evaluate sin π/5 , sin 2π/5 and cos π/5 .

10.2 Let z = cos θ + i sin θ.

Then zn = cos(nθ) + i sin(nθ) for all n ∈ N (by de Moivre) and z−n = cos(nθ) − i sin(nθ).

(a) Show that 2 cos(nθ) = zn + z−n and 2i sin(nθ) = zn − z−n.
(2)

(b) Show that 2n cosn θ = (z + 1 )n and (2i)n sinn θ = (z − 1 )n.
(2)



(c) Use (b) to express sin7 θ in terms of multiple angles.
(6)

(d) Express cos3 θ sin4 θ in terms of multiple angles.
(6)

(e) Eliminate θ from the equations 4x = cos(3θ) + 3 cos θ; 4y = 3 sin θ − sin(3θ).
(5)

[30]
1
Expert's answer
2020-05-05T20:47:50-0400

10.1

Let "\theta=\frac{\pi}{10}=18^0" then "5\theta=\frac{\pi}{2}"

"\cos3\theta=\sin2\theta\\
\cos54^0=\sin36^0=\sin(90^0-54^0)\\
4\cos^3\theta-3\cos\theta=2\sin\theta\cos\theta\\
4\cos^2\theta-3=2\sin\theta\\
4(1-\sin^2\theta)-3=2\sin\theta\\
4\sin^2\theta+2\sin\theta-1=0\\
\sin\theta=\frac{\sqrt5-1}{4}\\
\cos2\theta=\cos\frac{\pi}{5}=1-2\sin^2\theta=\frac{\sqrt5+1}{4}\\
\sin\frac{\pi}{5}=\sqrt{1-\cos^2\theta}=\sqrt{1-(\frac{\sqrt5+1}{4})^2}=\frac{\sqrt{10-2\sqrt5}}{4}\\
\sin\frac{2\pi}{5}=2\sin\frac{\pi}{5}\cos\frac{\pi}{5}=\frac{\sqrt{10-2\sqrt5}}{4}\frac{\sqrt5+1}{2}=\\
=\frac{\sqrt{10+2\sqrt5}}{4}"

10.2

"z=\cos\theta+i\sin\theta\\
z^n=\cos n\theta+i\sin n\theta\\
z^{-n}=\frac{1}{\cos n\theta+i\sin n\theta}=\cos n\theta-i\sin n\theta\\
n\in N"

(a)

"z^n+z^{-n}=2\cos n\theta\\
z^n-z^{-n}=2i\sin n\theta"

(b)

"z=\cos\theta+i\sin\theta\\
(z+1)^n=(\cos\theta+i\sin\theta+1)^n=\\
=(2\cos^2\frac{\theta}{2}+i2\sin\frac{\theta}{2}\cos\frac{\theta}{2})^n=\\
=2^n\cos^n\frac{\theta}{2}(\cos\frac{\theta}{2}+i\sin\frac{\theta}{2})^n=\\
=2^n\cos^n\frac{\theta}{2}(\cos\frac{n\theta}{2}+i\sin\frac{n\theta}{2})=\\
=2^n\cos^n\frac{\theta}{2}z^{\frac{n}{2}}\\
2^n\cos^n\frac{\theta}{2}z^{\frac{n}{2}}=2^n\cos^ n\theta\\"


"(z-1)^n=(\cos\theta+i\sin\theta-1)^n=\\
=(2\sin^2\frac{\theta}{2}+i2\sin\frac{\theta}{2}\cos\frac{\theta}{2})^n=\\
=2^n\sin^n\frac{\theta}{2}(\sin\frac{\theta}{2}+i\cos\frac{\theta}{2})^n=\\
=2^n\sin^n\frac{\theta}{2}(-i(\cos\frac{\theta}{2}+i\sin\frac{\theta}{2}))^n=\\
=(2i)^n\sin^n\frac{\theta}{2}(\cos\frac{n\theta}{2}+i\sin\frac{n\theta}{2})=\\
=(2i)^n\sin^n\frac{\theta}{2}z^{\frac{n}{2}}\\
(2i)^n\sin^n\frac{\theta}{2}z^{\frac{n}{2}}=(2i)^n\sin^ n\theta\\"

(c)

"\sin\theta=\frac{z-z^{-1}}{2i}\\
\sin^7\theta=\frac{(z-z^{-1})^7}{2i}=\\
=\frac{1}{2i}(z^7-7z^6z^{-1}+21z^5z^{-2}-35z^4z^{-3}+\\+35z^3z^{-4}
-21z^2z^{-5}+7zz^{-6}-z^{-7})=\\
=\frac{1}{2i}(2i\sin7\theta-7\cdot(2i)\sin5\theta+21\cdot(2i)\sin3\theta-\\
-35(2i)\sin\theta)=\\
=sin7\theta-7\sin5\theta+21\sin3\theta-35\sin\theta"


(d)

"\cos3\theta=\frac{z^3+z^{-3}}{2}=\frac{(z+z^{-1})(z^2-zz^{-1}+z^{-2})}{2}=\\
=\frac{2\cos\theta(2\cos2\theta-1)}{2}=\cos\theta(2\cos2\theta-1)\\
\sin4\theta=\frac{z^4-z^{-4}}{2i}=\frac{(z^2+z^{-2})(z^2-z^{-2})}{2i}=\frac{2\cos2\theta2i\sin2\theta}{2i}=\\
=2\cos2\theta\sin2\theta"

(e)

"\cos3\theta=\frac{z^3+z^{-3}}{2}=\frac{(z+z^{-1})(z^2-zz^{-1}+z^{-2})}{2}=\\
=\frac{2\cos\theta(2\cos2\theta-1)}{2}=\cos\theta(4\cos^2\theta-3)=\\
=4\cos^3\theta-3\cos\theta\\
4x=\cos3\theta+3\cos\theta\\
4x=4\cos^3\theta-3\cos\theta+3\cos\theta\\
x=\cos^3\theta"


"\sin3\theta=\frac{z^3-z^{-3}}{2i}=\frac{(z-z^{-1})(z^2+zz^{-1}+z^{-2})}{2i}=\\
=\frac{2i\sin\theta(2\cos2\theta+1)}{2i}=\sin\theta(3-4\sin^2\theta)=\\
=3\sin\theta-4\sin^3\theta\\
4y=3\sin\theta-\sin3\theta\\
4y=3\sin\theta-3\sin\theta+4\sin^3\theta\\
y=\sin^3\theta"